OpenStudy (anonymous):

Points (5, 3) and (6, 2) lie on line k. What is the slope of the line that is parallel to k? would the answer be -1?

OpenStudy (anonymous):

What's the slope of the line between (5, 3) and (6, 2)? m = (2 - 3)/(6 - 5) = -1/1 = -1 Since line k has a slope of -1, then any line parallel to it also has a slope of -1.